# Ring formation from black hole superradiance through repeated particle production on bound orbits

### v1

@@ Acknowledgements.
This work is supported in part by the National Key Research and Development Program of China Grants No. 2020YFC2201501 and No. 2021YFC2203002, in part by the National Natural Science Foundation of China Grants No. 12475067, No. 12347103, No. 12235019, No. 12075297 and No. 12147103, in part by the Science Research Grants from the China Manned Space Project with NO. CMS-CSST-2021-B01, in part by the Fundamental Research Funds for the Central Universities.
@@ Appendix A Analysis of Additional Interaction Terms
Linear $\chi$ interactions: $\phi\chi$, $\phi^{2}\chi$, and $\phi^{3}\chi$
These couplings induce forced, rather than parametric, resonance. For instance, For example, with a background oscillation of the $\phi$ field—$\phi(t)=\Phi_{0}\cos(\mu_{\phi}t)$—the ${\phi}^{2}\chi$ interaction produces an inhomogeneous equation of motion
@@
Cubic $\chi$ interaction term $\phi\chi^{3}$
For the non-linear interaction $V_{\mathrm{int}}\propto\kappa\phi\chi^{3}$, the equation of motion for $\chi$ under the oscillating background of the $\phi$ cloud becomes:
@@
Quartic interaction $\phi^{2}\chi^{2}$
The quartic coupling $V_{\rm int}=\frac{1}{2}g^{2}\phi^{2}\chi^{2}$ leads to parametric resonance. The equation of motion for a $\chi_{k}$ mode is a Mathieu-type equation:
@@
This implies that the interaction energy term must be much larger than the particle’s bare mass squared, $g^{2}\Phi_{0}^{2}\gg\mu_{\chi}^{2}$. In this regime, the dynamics is dominated by the non-gravitational force proportional to $g^{2}\nabla(\phi^{2})$. Our numerical simulations and stability analysis (see Appendix C) of the particle trajectories both confirm that this dominant force renders the orbits unstable and rapidly accelerates the produced $\chi$ particles to relativistic energies. This process is illustrated in a benchmark simulation in Fig. 4 and Fig. 5. The particle is quickly ejected on an escaping trajectory, and its Lorentz factor, after an initial period of rapid oscillation, grows to a large, constant value, confirming the particle becomes ultra-relativistic. This result is consistent with the findings of Ref. Chen et al. (2025), which demonstrated through a similar analysis that a strong gradient force from a scalar cloud can accelerate fermions to energies on the order of $g\Phi_{0}$. The final state is therefore an escaping flux rather than a stable, bound ring.
@@ Appendix B WKB and Weak-Field Approximation
$-2m_{\mathrm{eff}}\,k^{\alpha}\nabla_{\alpha}m_{\mathrm{eff}}\,.$ (36)
@@
The term on the right-hand side acts as a non-gravitational force, arising from the gradient of the $\phi$ field, which deflects the particle from a standard geodesic path, consistent with similar analyses for fermions in Ref. Chen et al. (2025).
@@ Appendix C Orbital Stability Analysis
In this appendix, we provide a stability analysis for the quasi-circular particle orbits that are essential for ring formation. Our starting point is the full, non-linear equation of motion Eq.(19), which is derived from the Lagrangian in the weak-field limit. We present here in its spatial component form:
@@
To analyze the orbit stability, we linearize these equations around a circular reference path in the equatorial plane, defined by $r(t)=2r_{0}$, $\theta(t)=\pi/2$ and $\varphi=Ct$, where $C$ is some constant angular velocity. We introduce small perturbations to this state: ${\mathbf{r}}(t)=\mathbf{r}_{0}(t)+\delta\mathbf{r}(t)$. After a lengthy but straightforward process of linearization and converting to dimensionless coordinates $x=r/(2r_{0})$ and $z=\Omega t$, we obtain a set of coupled linear ordinary differential equations for the perturbations. Notably, the vertical motion $\delta\theta$ decouples from the equatorial plane motion ($\delta x$, $\delta\varphi$). The linearized equations are:
@@ Appendix D Derivation of the Averaged Backreaction Term
The analysis begins with the formal expression for the expectation value of the squared field operator. In the presence of particle production, this can be expressed in terms of the Bogoliubov coefficients, $\alpha_{k}$ and $\beta_{k}$, for each momentum mode $k$. The expectation value contains a contribution from the produced particles (with occupation number $n_{k}=|\beta_{k}|^{2}$) and an oscillatory term Kofman et al. (1997):
@@
[…]

### v2

@@ Acknowledgements.
This work is supported in part by the National Key Research and Development Program of China Grants No. 2020YFC2201501 and No. 2021YFC2203002, in part by the National Natural Science Foundation of China Grants No. 12475067, No. 12347103, No. 12235019, No. 12075297, No. 12147103 and No.12588101, in part by the Science Research Grants from the China Manned Space Project with NO. CMS-CSST-2021-B01, in part by the Fundamental Research Funds for the Central Universities.
@@ Appendix A Analysis of Additional Interaction Terms
[…]
@@
Cubic $\chi$ interaction term $\phi\chi^{3}$—For the nonlinear interaction $V_{\mathrm{int}}\propto\kappa\phi\chi^{3}$, the equation of motion for $\chi$ under the oscillating background of the $\phi$ cloud becomes
@@
Quartic interaction $\phi^{2}\chi^{2}$—The quartic coupling $V_{\rm int}=\frac{1}{2}g^{2}\phi^{2}\chi^{2}$ leads to parametric resonance. The equation of motion for a $\chi_{k}$ mode is a Mathieu-type equation:
@@
This implies that the interaction energy term must be much larger than the particle’s bare mass squared, $g^{2}\Phi_{0}^{2}\gg\mu_{\chi}^{2}$. In this regime, the dynamics is dominated by the nongravitational force proportional to $g^{2}\nabla(\phi^{2})$. Our numerical simulations and stability analysis (see Appendix C) of the particle trajectories both confirm that this dominant force renders the orbits unstable and rapidly accelerates the produced $\chi$ particles to relativistic energies. This process is illustrated in a benchmark simulation in Fig. 4 and Fig. 5. The particle is quickly ejected on an escaping trajectory, and its Lorentz factor, after an initial period of rapid oscillation, grows to a large, constant value, confirming the particle becomes ultra-relativistic. This result is consistent with the findings of Ref. Chen:2023vkq, which demonstrated through a similar analysis that a strong gradient force from a scalar cloud can accelerate fermions to energies on the order of $g\Phi_{0}$. The final state is therefore an escaping flux rather than a stable, bound ring.
@@ Appendix B WKB and Weak-Field Approximation
$2m_{\mathrm{eff}}\,k^{\alpha}\nabla_{\alpha}m_{\mathrm{eff}}\,.$ (36)
@@
The term on the right-hand side acts as a nongravitational force, arising from the gradient of the $\phi$ field, which deflects the particle from a standard geodesic path, consistent with similar analyses for fermions in Ref. Chen:2023vkq.
@@ Appendix C Orbital Stability Analysis
In this Appendix, we provide a stability analysis for the quasicircular particle orbits that are essential for ring formation. Our starting point is the full, nonlinear equation of motion Eq.(19), which is derived from the Lagrangian in the weak-field limit. We present here in its spatial component form:
@@
To analyze the orbit stability, we linearize these equations around a circular reference path in the equatorial plane, defined by $r(t)=2r_{0}$, $\theta(t)=\pi/2$ and $\varphi=Ct$, where $C$ is some constant angular velocity. We introduce small perturbations to this state: ${\mathbf{r}}(t)=\mathbf{r}_{0}(t)+\delta\mathbf{r}(t)$. After a lengthy but straightforward process of linearization and converting to dimensionless coordinates $x=r/(2r_{0})$ and $z=\Omega t$, we obtain a set of coupled linear ordinary differential equations for the perturbations. Notably, the vertical motion $\delta\theta$ decouples from the equatorial plane motion ($\delta x$, $\delta\varphi$). The linearized equations are
@@ Appendix D Derivation of the Averaged Backreaction Term
The analysis begins with the formal expression for the expectation value of the squared field operator. In the presence of particle production, this can be expressed in terms of the Bogoliubov coefficients, $\alpha_{k}$ and $\beta_{k}$, for each momentum mode $k$. The expectation value contains a contribution from the produced particles (with occupation number $n_{k}=|\beta_{k}|^{2}$) and an oscillatory term Kofman:1997yn:
@@
where $C$ is some constant. Consequently, the phase term $e^{-2i\int\omega_{k}dt^{\prime}}$ oscillates at a very high frequency compared to the evolution of the background field $\phi(t)$ itself, which occurs at the low frequency $\mu_{\phi}$. When $\langle\chi^{2}\rangle$ acts as a source in the equation of motion for $\phi$, the contribution from this rapidly oscillating term is negligible Kofman:1997yn. Therefore, for modeling the backreaction on the evolution of the $\phi$ cloud, the expectation value is dominated by the particle number term: $\overline{\langle\chi^{2}\rangle}\approx\int\frac{d^{3}k}{(2\pi)^{3}}\frac{n_{k}}{\omega_{k}(t)}$.
By defining the total number density as $n_{\chi}=\int\frac{d^{3}k}{(2\pi)^{3}}n_{k}$, we arrive at the simplified expression used in the main text: